19F Spin-rotation Constants and Chemical Shifts of Unstable Molecules from their Microwave Spectra

\underline{Bethany Gatehouse}, Holger S. P. Müller^{a}, and Michael C. L. Gerry, University of British Columbia, Department of Chemistry, 2036 Main Mall, Vancouver, B.C., Canada, V6T 1Z1.

The inherently high resolution of the pulsed molecular beam-microwave Fourier transform (MWFT) technique allows spin-rotation coupling constants to be routinely determined from small hyperfine splittings in rotational transitions. The relationship between the spin-rotation coupling constants and the chemical shielding of a given nucleus has been well established^{b}; these constants can be used to calculate the individual components of the chemical shielding tensor as well as the average chemical shift. Thus, shielding data for unstable and semi-stable molecules, which are relatively easily studied by MWFT, can be obtained. Several ^{19}F containing species have been prepared using an electric discharge. Their microwave spectra have been measured in the 4-26 GHz frequency range and the determined diagonal components of the nuclear spinrotation tensors have been used to calculate the chemical shielding tensors. One such molecule that is of particular interest is sulphur difluoride. Two of its spin-rotation coupling constants were found to be negative. Negative fluorine spin-rotation coupling constants are very unusual, and have been seen before only for the interhalogens^{c} ClF, BrF, and IF. These findings are consistent with previous ab initio calculations^{d} and an NMR experiment^{e}.

apresent address: Jet Propulsion Laboratory, California Institute of Technology, Pasadena, California, 91109

bW. H. Flygare, J. Chem. Phys., 41, 793, (1964)

cH. S. P. Müller and M. C. L. Gerry, J. Chem. Phys., 103, 577, (1995); and references therein

dM. Schindler, J. Chem. Phys., 88, 7638, (1988)

eW. Gombler, J. Schaebs, and H. Willner, Inorg. Chem., 29, 2697, (1990)